Session: Mar 2007 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K. Borkowski at Tue Mar 20 09:25:05 2007) Minor failures : RECE
Observing only in one channel (see notes for F07M1 and N07M1). As last night CL experiment indicates, Tsys is high but not that bad I reported earlier using FS results obtained with old rxg data.

Effelsberg (A. Polatidis at Tue Mar 20 09:41:54 2007) Success :
No known problems

Medicina (G.Maccaferri at Tue Mar 20 09:51:52 2007) Minor failures : RFI

Hartebeesthoek (Jonathan Quick at Tue Mar 20 12:58:17 2007) Minor failures : FS
Local control system glitch cause us to miss scan No0101. No other known problems.

Jodrell_Bank (Alastair Gunn at Tue Mar 20 16:06:30 2007) Success :
The MkII telescope was used (Jb2). No known problems.

Cambridge (Alastair Gunn at Tue Mar 20 16:07:21 2007) Minor failures : FS
FS computer failed at about 05:03 UT. Managed to get system running for 06:16 UT scan. No other known problems.

Westerbork (Tony Foley at Wed Mar 21 09:08:10 2007) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. Looks fine.

Noto (P. Cassaro at Sun Mar 25 23:29:04 2007) Success :